# CCFQA: A Benchmark for Cross-Lingual and Cross-Modal Speech and Text Factuality Evaluation

### v1

@@ Abstract
As Large Language Models (LLMs) are increasingly popularized in the multilingual world, ensuring hallucination-free factuality becomes markedly crucial. However, existing benchmarks for evaluating the reliability of Multimodal Large Language Models (MLLMs) predominantly focus on textual or visual modalities with a primary emphasis on English, which creates a gap in evaluation when processing multilingual input, especially in speech. To bridge this gap, we propose a novel Cross-lingual and Cross-modal Factuality benchmark (CCFQA). Specifically, the CCFQA benchmark contains parallel speech-text factual questions across 8 languages, designed to systematically evaluate MLLMs’ cross-lingual and cross-modal factuality capabilities. Our experimental results demonstrate that current MLLMs still face substantial challenges on the CCFQA benchmark. Furthermore, we propose a few-shot transfer learning strategy that effectively transfers the Question Answering (QA) capabilities of LLMs in English to multilingual Spoken Question Answering (SQA) tasks, achieving competitive performance with GPT-4o-mini-Audio using just 5-shot training. We release CCFQA as a foundational research resource to promote the development of MLLMs with more robust and reliable speech understanding capabilities. [^1] Code and Dataset: https://github.com/yxduir/ccfqa
@@ Introduction
Large Language Models (LLMs) have achieved significant progress in recent years, driving remarkable advancements across numerous fields and applications. However, the inherent issue of hallucinations (Rawte, Sheth, and Das 2023) in LLMs remains unresolved, especially when faced with multilingual and multimodal inputs. With the popularization of Multimodal Large Language Models (MLLMs) (Chen et al. 2024a), an increasing number of rich multilingual multimedia scenarios have emerged, such as spoken dialogue and digital humans, where the hallucination problem becomes particularly pronounced. This issue is further exacerbated in cross-lingual and cross-modal settings. As shown in Figure 1, even the GPT series models struggle to mitigate hallucinations in these complex scenarios. That is, MLLMs may yield inconsistent answers when the same factual question is asked to be answered in different languages or presented via different input modalities, leading to poor user experience and limiting the global adoption of LLMs.
Factuality benchmarks, such as SimpleQA (Wei et al. 2024), have gained increasing attention as effective tools for assessing hallucination in LLMs. These benchmarks use fact-based QA tasks to objectively evaluate model accuracy and reliability. However, most existing benchmarks (Zhang et al. 2025) focus on textual or visual inputs and are primarily designed for English, lacking coverage of multilingual speech scenarios. As shown in Table 1, a comprehensive benchmark for evaluating multilingual speech settings is still missing.
@@
TruthfulQA (Lin, Hilton, and Evans 2022) English 817 Acc
HaluEval (Li et al. 2023a) English 5,000 Acc
SimpleQA (Wei et al. 2024) English 4,326 LLM
Chinese SimpleQA (He et al. 2024) Chinese 3,000 LLM
KoLasSimpleQA (Jiang et al. 2025) 9 2,147 LLM
SD-QA (Faisal et al. 2021) 5 11,109 F1
CompA (Ghosh et al. 2024) English 600 Acc
VoiceBench (Chen et al. 2024b) English 5,783 Acc / LLM mix
SpeechIQ (Wan et al. 2025) English 800 LLM
@@
Fact-based question answering is a vital research area for evaluating LLMs. TruthfulQA (Lin, Hilton, and Evans 2022), known as the first fact-based benchmark in the era of LLMs, assesses truthfulness by measuring how LLMs handle questions involving common human misconceptions. Its goal is to prevent models from generating imitative falsehoods learned from their training data. HaluEval (Li et al. 2023a) addresses ”hallucinations,” which are coherent but factually incorrect texts, by using human-annotated samples. SimpleQA (Wei et al. 2024) employs adversarially collected, short, fact-seeking questions with single answers to test a model’s factual recall and its “knowledge of its own knowledge.” Extending fact-based evaluation to other languages, Chinese SimpleQA (He et al. 2024) is the first comprehensive Chinese benchmark with high-quality answers, mainly focusing on evaluating Chinese-centered LLMs. Finally, KoLasSimpleQA (Jiang et al. 2025) is the first multilingual fact-based benchmark, assessing nine languages on both general and language-specific knowledge.
@@
Evaluating Spoken Question Answering (SQA) is an evolving field. SD-QA (Faisal et al. 2021) is a typical work addressing dialectal variations with a dataset of over 68,000 spoken prompts, enabling real-world performance and fairness evaluations. CompA (Ghosh et al. 2024), while not a direct benchmark, offers insights into multisensory fusion relevant for designing multi-modal SQA systems. VoiceBench (Chen et al. 2024b) provides a comprehensive benchmark for LLM-based voice assistants, assessing general knowledge, instruction-following, and safety under realistic audio conditions. SpeechIQ (Wan et al. 2025) introduces a cognition-inspired pipeline, SIQ, which evaluates models across three levels of Bloom’s Taxonomy, providing a holistic assessment of a model’s understanding. All of these benchmarks focus on a single voice modality and use non-open-ended answers, highlighting a common limitation in current SQA evaluation methods.
@@
To construct the benchmark, we leverage text-based QA datasets: MKQA (Longpre, Lu, and Daiber 2021) and MOOCCubeX (Yu et al. 2021). MKQA provides a wealth of open-domain questions, with the coverage of general knowledge topics such as movies, music, and sports. MOOCCubeX offers questions rooted in educational contexts, covering a wide array of academic subjects.
@@
Figure 3: The Question Categories and Benchmark Statistics in the CCFQA Benchmark.
@@
To enforce a final layer of quality control on the audio data, we implement a systematic speech refinement loop. We apply audio augmentation to enhance samples with low volume levels. Additionally, we employ Whisper-large-v3 (Radford et al. 2023), an Automatic Speech Recognition (ASR) model, for a comprehensive quality check. The ASR system transcribes every recorded audio file. We then calculate the Word Error Rate (WER) and Character Error Rate (CER) by comparing the ASR-generated transcripts with the original ground-truth text.
@@
Figure 4: The Architecture of LLM-SQA. LLM-SQA consists of a speech encoder, speech adapter, and LLM. A curriculum learning strategy sequentially trains the ASR, SRT, and SQA tasks. We train the <|qa|> task in English, and for other languages, cross-lingual knowledge transfer can be achieved with only 5-shot trainings, using English as a bridge.
@@
Following the design of LLM-SRT (Du et al. 2025), we adopt a sequential curriculum learning strategy that trains the model on three tasks in order: (1) Automatic Speech Recognition (ASR), (2) Speech Recognition and Translation (SRT), and (3) Spoken Question Answering (SQA). To help the model distinguish between different tasks, we design explicit task instructions that ensure clear separation between instructions and predicted outputs, as shown in Figure 4. The ASR and SRT tasks are trained on the FLEURS dataset (Conneau et al. 2023).
@@
Our MLLM is pre-trained from scratch based on GemmaX2-9B (Cui et al. 2025), with added support for Cantonese (Hong Kong SAR). Our model supports ASR, Speech-to-Text Translation (S2TT), and SQA tasks together within a single query.
@@
Our MLLM includes a frozen speech encoder, a trainable adapter composed of a Q-Former (Li et al. 2023b) with 80 queries of dimension 768 and an MLP, and an LLM, as detailed in Table 3.
@@
Experiments run for one week on four A100 (80GB) GPUs. We use AdamW (Loshchilov 2017) with a peak learning rate of $1\times 10^{-4}$, 1000-step warm-up, and linear decay thereafter.
@@
We compare five MLLMs: GPT-4o-mini (OpenAI 2023), Phi-4-Multimodal (Abouelenin et al. 2025), Qwen2-Audio (Chu et al. 2024), and the Qwen2.5-Omni series (Xu et al. 2025).
@@
Table 4: F1 and LLM-based accuracy of MLLMs on the CCFQA benchmark, including cross-lingual (QA $\rightarrow$ XQA / SQA $\rightarrow$ XSQA) and cross-modal (QA $\rightarrow$ SQA / XQA $\rightarrow$ XSQA) settings. Performance degradation is observed in both cross-lingual and cross-modal scenarios. Full results are provided in Tables 10 and $1$ in the Appendix.
@@
Special Tokens for MLLMs.
@@ Appendix A Appendix
Specifically, Gemma3-27B (Team et al. 2025) is selected as the judge LLM, which, given a question, an answer, the language direction, and a response, judges whether the response is correct by answering ”yes” or ”no.” For the same task, all models are evaluated using identical prompts and consistent sampling parameters set via vLLM (Kwon et al. 2023). We selected the LLM as our evaluation model based on the following key factors:

### v2

@@ Abstract
As Large Language Models (LLMs) are increasingly popularized in the multilingual world, ensuring hallucination-free factuality becomes markedly crucial. However, existing benchmarks for evaluating the reliability of Multimodal Large Language Models (MLLMs) predominantly focus on textual or visual modalities with a primary emphasis on English, which creates a gap in evaluation when processing multilingual input, especially in speech. To bridge this gap, we propose a novel Cross-lingual and Cross-modal Factuality benchmark (CCFQA). Specifically, the CCFQA benchmark contains parallel speech-text factual questions across 8 languages, designed to systematically evaluate MLLMs’ cross-lingual and cross-modal factuality capabilities. Our experimental results demonstrate that current MLLMs still face substantial challenges on the CCFQA benchmark. Furthermore, we propose a few-shot transfer learning strategy that effectively transfers the Question Answering (QA) capabilities of LLMs in English to multilingual Spoken Question Answering (SQA) tasks, achieving competitive performance with GPT-4o-mini-Audio using just 5-shot training. We release CCFQA as a foundational research resource to promote the development of MLLMs with more robust and reliable speech understanding capabilities.
@@ Introduction
Large Language Models (LLMs) have achieved significant progress in recent years, driving remarkable advancements across numerous fields and applications. The popularization of Multimodal Large Language Models (MLLMs) (li_unimoe; lmeye) has amplified the hallucination (rawte2023survey) problem, particularly in rich multilingual scenarios. This issue is further exacerbated in cross-lingual and cross-modal settings. As shown in Figure 1, even the GPT series models struggle to mitigate hallucinations in these complex scenarios. That is, MLLMs may yield inconsistent answers when the same factual question is asked to be answered in different languages or presented via different input modalities.
Factuality benchmarks, such as SimpleQA (Wei2024MeasuringSF), have gained increasing attention as effective tools for assessing hallucination in LLMs. These benchmarks use fact-based QA tasks to objectively evaluate model accuracy and reliability. However, most existing benchmarks (zhang-etal-2025-cchall) focus on textual or visual inputs and are primarily designed for English, lacking coverage of multilingual speech scenarios. As shown in Table 1, a comprehensive benchmark for evaluating multilingual speech settings is still missing.
@@
TruthfulQA (lin2022truthfulqa) English 817 Acc
HaluEval (HaluEval) English 5,000 Acc
SimpleQA (Wei2024MeasuringSF) English 4,326 LLM
Chinese SimpleQA (he2024chinesesimpleqachinesefactuality) Chinese 3,000 LLM
KoLasSimpleQA (jiang2025evaluating) 9 2,147 LLM
SD-QA (faisal2021sd) 5 11,109 F1
CompA (ghosh2024compa) English 600 Acc
VoiceBench (chen2024voicebench) English 5,783 Acc / LLM mix
SpeechIQ (wan2025speechiq) English 800 LLM
@@
Recently, a series of evaluation benchmarks (liu-etal-2025-projecteval) have emerged in the LLM field, with factuality benchmarks being one of them. Fact-based question answering is a vital research area for evaluating LLMs. TruthfulQA (lin2022truthfulqa), known as the first fact-based benchmark in the era of LLMs, assesses truthfulness by measuring how LLMs handle questions involving common human misconceptions. Its goal is to prevent models from generating imitative falsehoods learned from their training data. HaluEval (HaluEval) addresses ”hallucinations,” which are coherent but factually incorrect texts, by using human-annotated samples. SimpleQA (Wei2024MeasuringSF) employs adversarially collected, short, fact-seeking questions with single answers to test a model’s factual recall. Extending fact-based evaluation to other languages, Chinese SimpleQA (he2024chinesesimpleqachinesefactuality) is the first comprehensive Chinese benchmark with high-quality answers, mainly focusing on evaluating Chinese-centered LLMs. Finally, KoLasSimpleQA (jiang2025evaluating) is the first multilingual fact-based benchmark, assessing nine languages on both general and language-specific knowledge.
@@
Evaluating Spoken Question Answering (SQA) is an evolving field. SD-QA (faisal2021sd) is a typical work addressing dialectal variations with a dataset of over 68,000 spoken prompts, enabling real-world performance and fairness evaluations. CompA (ghosh2024compa), while not a direct benchmark, offers insights into multisensory fusion relevant for designing multi-modal SQA systems. VoiceBench (chen2024voicebench) provides a comprehensive benchmark for LLM-based voice assistants, assessing general knowledge, instruction-following, and safety under realistic audio conditions. SpeechIQ (wan2025speechiq) introduces a cognition-inspired pipeline, SIQ, which evaluates models across three levels of Bloom’s Taxonomy, providing a holistic assessment of a model’s understanding. All of these benchmarks focus on a single voice modality and use non-open-ended answers, highlighting a common limitation in current SQA evaluation methods.
@@
To construct the benchmark, we leverage text-based QA datasets: MKQA (longpre2021mkqa) and MOOCCubeX (yu2021mooccubex). MKQA provides a wealth of open-domain questions, with the coverage of general knowledge topics such as movies, music, and sports. MOOCCubeX offers questions rooted in educational contexts, covering a wide array of academic subjects.
@@
To enforce a final layer of quality control on the audio data, we implement a systematic speech refinement loop. We apply audio augmentation to enhance samples with low volume levels. Additionally, we employ Whisper-large-v3 (whisper), an Automatic Speech Recognition (ASR) model, for a comprehensive quality check. The ASR system transcribes every recorded audio file. We then calculate the Word Error Rate (WER) and Character Error Rate (CER) by comparing the ASR-generated transcripts with the original ground-truth text.
@@
Figure 3: The Question Categories in the CCFQA.
Table 3: Benchmark Statistics. CCFQA supports the evaluation of QA, XQA, SQA, and XSQA for MLLMs. It features parallel speech-text question pairs across 8 languages. Due to the cross-lingual evaluation settings in XQA and XSQA, a full evaluation of one MLLM requires 128,000 individual requests.
@@
Figure 4: The Architecture of LLM-SQA.
@@
Following the design of LLM-SRT (du2025makingllmsbettermanytomany), we adopt a sequential curriculum learning strategy that trains the model on three tasks in order: (1) Automatic Speech Recognition (ASR), (2) Speech Recognition and Translation (SRT), and (3) Spoken Question Answering (SQA). To help the model distinguish between different tasks, we design explicit task instructions that ensure clear separation between instructions and predicted outputs, as shown in Figure 4. The ASR and SRT tasks are trained on the FLEURS dataset (conneau2023fleurs).
@@
Our MLLM is pre-trained from scratch based on GemmaX2-9B (cui2025multilingual), with added support for Cantonese (Hong Kong SAR). Our model supports ASR, Speech-to-Text Translation (S2TT), and SQA tasks together within a single query.
Task Instruction Prediction
ASR <|eng|> what is the largest city in the uk
<|fra|> quelle est la plus grande ville du royaume-uni
<|spa|> cual es la ciudad mas grande de uk
SRT <|eng|><|fra|> what is the largest city in the uk<|eng|><|fra|>
quelle est la plus grande ville du royaume-uni
<|spa|><|eng|> cual es la ciudad mas grande de uk<|spa|><|eng|>
what is the largest city in the uk
SQA <|qa|> what is the largest city in the uk
<|qa|>London
<|spa|><|eng|> cual es la ciudad mas grande de uk<|spa|><|eng|>
<|qa|> what is the largest city in the uk
<|qa|>London
<|spa|><|eng|> cual es la ciudad mas grande de uk<|spa|><|eng|>
<|qa|><|fra|> what is the largest city in the uk
<|qa|>London<|fra|>Londres
Table 4: Instruction Design. LLM-SQA consists of a speech encoder, speech adapter, and LLM. A curriculum learning strategy sequentially trains the ASR, SRT, and SQA tasks. We train the <|qa|> task in English, and for other languages, cross-lingual knowledge transfer can be achieved with only 5-shot trainings, using English as a bridge.
@@
Our MLLM includes a frozen speech encoder, a trainable adapter composed of a Q-Former with 80 queries of dimension 768 and an MLP, and an LLM, as detailed in Table 5.
@@
Experiments run for one week on four A100 (80GB) GPUs. We use AdamW with a peak learning rate of $1\times 10^{-4}$, 1000-step warm-up, and linear decay thereafter.
@@
We compare five MLLMs: GPT-4o-mini (achiam2023gpt), Phi-4-Multimodal (abouelenin2025phi), Qwen2-Audio (chu2024qwen2), and the Qwen2.5-Omni series (xu2025qwen2).
@@
Table 6: F1 and LLM-based accuracy of MLLMs on the CCFQA benchmark, including cross-lingual (QA $\rightarrow$ XQA / SQA $\rightarrow$ XSQA) and cross-modal (QA $\rightarrow$ SQA / XQA $\rightarrow$ XSQA) settings. Performance degradation is observed in both cross-lingual and cross-modal scenarios.
@@
Language Tags for MLLMs.
@@ Conclusion
Acknowledgements
The research in this article is supported by the National Science Foundation of China (U22B2059, 62276083, 62506182), the 5G Application Innovation Joint Research Institute’s Project (A003), the Major Key Project of PCL (Grant No. PCL2025A12, PCL2025A03), the National Science and Technology Major Program (Grant No. 2024ZD01NL00101).
@@ Appendix A Appendix
Specifically, Gemma3-27B (team2025gemma) is selected as the judge LLM, which, given a question, an answer, the language direction, and a response, judges whether the response is correct by answering ”yes” or ”no.” For the same task, all models are evaluated using identical prompts and consistent sampling parameters set via vLLM (kwon2023efficient). We selected the LLM as our evaluation model based on the following key factors: